At each vertex of the square (0, 0, 0), (3, 0, 0), (3, 3, 0), (0, 3, 0) there is a point charge of 4.0 (C). With center at point (5, 5, 0) there is a sphere of radius R = 5.65m how much is the value of ∯ E.ds when integrating over the entire surface of the sphere? Note: the coordinates of the points are given in meters.
Answer #1

R = 5.65 m 9,0 15,,0) 10 A $ (3,030) from Gausss law we know that : | 88. EdB = 9 total where, a total = Total Change in clo80= 8.85 X10-12 c/um. Sor q total = (4+4+4+4)C q total = 16 c. $ Eids = 16 e 8:85x10-12 X/ V-m 2 m-n 21 ks = p.28 - - $7:43 =